4. Tuttle Buildings Inc. has decided to go public by selling $5,000,000 of new common stock. Its investment bankers agre versus their normal 10%) in exchange for a 200,000 shares at $5.00 per share. The investment bankers expect to exercise the ed to take a smaller fee now (6% of gross proceeds 1-year option to purchase an additional 200,000 shares in exactly one year, when the stock price is r share. However, there is a chance that the stock price will option and purchase the forecasted to be $6.50 pe actually be $12.00 per share one year from now. If the $12 price occurs, the present value of the entire underwriting compensation be? Assume that investment bankers required return on such ar what would the rangements is 15%, and ignore taxes.
Answer #1

The bankers fee = 6%*5000,000 = $300,000

Cash flow after 1 year = Number of shares * Profit on share

200000*(12-5)

=200000*7 = 1400000

PV of cash flows= 300000+ 1400000/1.15^1

= 300000+1217391

=$1517391